# Enhancing Transformers for Generalizable First-Order Logical Entailment

### v1

@@ Abstract
Transformers, as a fundamental deep learning architecture, have demonstrated remarkable capabilities in reasoning. This paper investigates the generalizable first-order logical reasoning ability of transformers with their parameterized knowledge and explores ways to improve it. The first-order reasoning capability of transformers is assessed through their ability to perform first-order logical entailment, which is quantitatively measured by their performance in answering knowledge graph queries. We establish connections between (1) two types of distribution shifts studied in out-of-distribution generalization and (2) the unseen knowledge and query settings discussed in the task of knowledge graph query answering, enabling a characterization of fine-grained generalizability. Results on our comprehensive dataset show that transformers outperform previous methods specifically designed for this task and provide detailed empirical evidence on the impact of input query syntax, token embedding, and transformer architectures on the reasoning capability of transformers. Interestingly, our findings reveal a mismatch between positional encoding and other design choices in transformer architectures employed in prior practices. This discovery motivates us to propose a more sophisticated, logic-aware architecture, TEGA, to enhance the capability for generalizable first-order logical entailment in transformers.
@@
Tianshi Zhengthanks:   Equal Contribution, Jiazheng Wang11footnotemark: 1, Zihao Wang, Jiaxin Bai, Hang Yin, Zheye Deng, Yangqiu Song, Jianxin Li Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China Department of Computer Science and Engineering, Beihang University, Beijing, China Department of Mathematical Sciences, Tsinghua University, Beijing, China tzhengad@connect.ust.hk
@@ 1 Introduction
Interestingly, our results demonstrated the superior performance and generalizability of relative positional encoding (RPE) over traditional absolute positional encoding (APE). However, previous studies have only proposed inductive biases within the APE setting, which we prove ineffective in the RPE setting. To fill this gap, we propose TEGA (Transformer Encoder with Guided Attention), a novel modeling methodology that facilitates effective reasoning with logic-aware guidance in self-attention. Our study shows that TEGA substantially improves the performance and generalizability upon transformers under the RPE setting. (§6) Our code and data can be found at https://anonymous.4open.science/r/TEGA.
@@ 2.2 Knowledge Graph Query Answering
3 Distribution Shifts in KG Query Answering
@@ 4 Evaluation Benchmark
Trans.+Disentangled PE Transformer 73.6 19.3 29.7 8.7 48.1 13.1
Trans.+Relative PE Transformer 68.0 18.7 49.8 13.4 57.4 15.6
@@ 4.3 General Benchmarking Results
We evaluate transformers against five existing methods for KG query answering, with details provided in Appendix B. Meanwhile, we compare three positional encoding settings in transformers: Absolute PE is the default sinusoidal positional encoding introduced in the original transformer Vaswani et al. (2023). Relative PE Shaw et al. (2018) applied a learnable matrix based on the relative position of tokens. Disentangled PE He et al. (2021) uses separate vectors to encode the content and position of each token, decoupling the contextual and positional information during self-attention.
The experimental results are presented in Table 2. Transformers marginally outperform all baselines in both ID (K) and OOD (K) in seen query types. Both transformers with absolute PE and disentangled PE have poor OOD generalizability in the query type dimension. In contrast, transformer with relative PE demonstrates much stronger generalizability, as it can better capture structural information from the distances between tokens and flexibly handle query sequences of unseen lengths. Overall, transformers with relative PE achieve the best performance and OOD generalizability on first-order logical entailment.
@@ 5 Transformer Architectures and First-order Reasoning Capability
In this section, we further investigate the dependencies between transformers’ performances and multiple design choices throughout the modeling process. Figure 2 depicts the three-stage pipeline of deriving the answer set $\mathcal{A}$ (output $Y$) of query $q$ (input $X$) using transformers. The impact of query syntax, token embeddings, and transformer architecture are presented in Section 5.1, Section 5.2, and Section 5.3, respectively, and summarized in Section 5.4.
@@ 5.1 Study on Query Syntax
Logical queries with the same semantics can be represented in different syntaxes with different formal languages. To investigate the impact of difference in formal languages, we conducted experiments on two formal languages: Lisp-like Syntax (Wang et al., 2021) represents query with fully parenthesized nested formula; EFO Syntax Yin et al. (2023b) represents query with all one-hop atomics connected with conjunctions and disjunctions in parallel. As the Lisp-like syntax has a limited representation scope, we used a subset of our benchmark that includes 13 seen query types and 12 unseen query types that are supported by both languages. The experimental results on the impact of formal language are presented in Table 3. In transformers with absolute PE, there is no difference between the results of the two formal languages. However, with relative PE, the EFO syntax achieves much better generalizability in the query type domain. This phenomenon can be explained by the differences in the structural features of the two languages: with a parallel structure, the distances (i.e., relative positions) between tokens with the same logical relationship are more consistent than those in a nested structure. As a result, relative PE can better learn the logical relationship between tokens and better generalize to unseen query types.
@@
Furthermore, even for the same query in EFO Syntax, there can be various permutations of one-hop atomics, with the total number of potential permutations growing factorially with the length of the query. Therefore, it is crucial for transformers to robustly handle queries with different permutations. To evaluate this property, we selected five query types (2p, 3p, ip, pi, 2in) and reversed their permutations while maintaining the same semantics in the testing data (see Appendix B). The experimental results on query permutation are presented in Table 4. We observed a significant performance drop after reversing the query in transformers with absolute PE. However, for transformers with relative PE, the performance before and after reversing the permutation remains consistent, demonstrating their robustness in handling queries with modified permutations.
Dataset KG Embedding All Queries
ID (K) OOD (K)
FB15k Random (learned) 40.7 26.0
TransE (frozen) Bordes et al. (2013) 38.4 24.8
DistMult (frozen) Yang et al. (2015) 42.6 26.4
@@
FB15k-237 Random (learned) 44.8 17.6
TransE (frozen) Bordes et al. (2013) 43.2 17.0
@@
NELL995 Random (learned) 57.4 15.6
TransE (frozen) Bordes et al. (2013) 57.0 15.0
[…]
@@
5.2 Study on Token Embeddings
Some methods in KG query answering Arakelyan et al. (2021); Wang et al. (2023b); Xu et al. (2023) utilized pre-trained KG embeddings from link predictors in their query encoding process, while other methods Ren and Leskovec (2020); Zhang et al. (2021); Bai et al. (2023b) randomly initialize the embedding matrix and jointly learn the KG embedding and model parameters from the training process. It’s still unclear that to what extent can pre-trained KG semantics can assist the performance of query answering models. To rigorously assess the impact of pre-trained semantics in token embeddings, we select three widely recognized pre-trained KG embeddings: TransE Bordes et al. (2013) models relationships by interpreting them as translations within the embedding space, DistMult Yang et al. (2015) employs a bilinear model wherein entities are represented as vectors and relations as diagonal matrices, ComplEx Trouillon et al. (2016) extends this framework into a complex vector space, allowing for more nuanced inter-entity interactions. In our experiment, all three KG embeddings are frozen during the training stage, compared to a baseline with randomly initialized and jointly learned embeddings.
The experimental results are presented in Table 5. Two stronger KGEs, ComplEx and DistMult, can effectively improve the performance of transformer. In contrast, TransE performance consistently lags behind the baseline in all three datasets. This discrepancy can be attributed to the embedding matrix learning process of the baseline model, which can be viewed as implicitly training a link predictor using a transformer architecture (e.g., KG-BERT Yao et al. (2019)). This modern approach may naturally outperform earlier models like TransE.
@@ 5.3 Study on Transformer Architecture
Some existing literature Kotnis et al. (2021); Liu et al. (2022); Xu et al. (2023) has explored the use and design of transformers in KG query answering, in which several inductive biases in transformer architecture were proposed. However, as all these methods focused on different aspects and were evaluated on different benchmarks, a clearer investigation is needed. We implemented two designs from previous works on top of transformers: Adjacency Matrices Masking: an inductive bias proposed in kgTransformer Liu et al. (2022), which aims to enhance reasoning performance by limiting the range of self-attention to one-hop neighboring nodes in the query graph. Directed Distance Encoding: another inductive bias, introduced in Query2Triple Xu et al. (2023), that facilitates information aggregation by injecting a directed distance message into the self-attention mechanism. The experimental results on existing transformer designs are presented in Table 6. Both inductive biases effectively improve generalizability in the query type dimension when using absolute positional encoding. In contrast, under relative positional encoding, these improvements are not observed, and the use of adjacency matrices masking even results in a significant decrease in performance. This decrease is attributed to the masking in self-attention, which restricts message passing to only nodes within one hop, thereby limiting the model’s ability to perform complex logical reasoning. Regarding directed distance encoding, the information about directed distances is mostly contained in the relative positions of the linearized sequence, thus the improvements seen with original absolute positional encoding are not sustained.
@@ 5.4 Summary of Empirical Findings
In terms of query syntax, we discovered that employing formal language with parallel structure can significantly enhance OOD generalizability in the query type dimension for transformer with relative PE. Moreover, we observed that transformers with relative PE can robustly handle queries with different permutations. For token embeddings, frozen pre-trained KGEs from stronger link predictors (e.g. ComplEx) outperform learned embeddings, whereas weaker ones (e.g. TransE) do not. Regarding transformer architecture, we found that the inductive biases proposed in previous approaches are only effective under absolute PE, but fail to improve upon transformer with relative PE, which we have shown to be the preferred setting. This mismatch between transformer design and positional encoding settings motivated us to design a methodology that effectively boosts the performance and generalizability of transformers with relative PE in first-order logical entailment.
@@ 7 Related Work
Transformers have exhibited remarkable performance in various forms of logical reasoning. In natural language inference (NLI) tasks Bowman et al. (2015); Williams et al. (2018), transformers analyze the logical relationship between a premise and a hypothesis with deductive reasoning over statements. Furthermore, transformers have been employed for inductive reasoning within rule-based systems Clark et al. (2020). Additionally, transformers have been applied to perform abductive reasoning in both natural language Bhagavatula et al. (2020) and formal language Bai et al. (2024) settings. Meanwhile, various techniques have been proposed to improve the logical reasoning capability of transformers for language understanding Chen (2023); Pan et al. (2023); Yuan et al. (2020). Though many inductive biases and empirical findings from natural language reasoning scenarios are not transferable to our setting due to the different nature of the task, some design choices may result in similar effects. For instance, the application of relative PE can improve the generalizability in both query answering and other natural language tasks Ontanon et al. (2022).
@@
Prior to this work, extensive research had been done on KG query answering with parameterized knowledge. Regarding modeling approaches, iterative neural query encoders Ren and Leskovec (2020); Zhang et al. (2021); Chen et al. (2022) design representations for entity sets and execute logical operators iteratively, following the computational graph. In addition, neural-symbolic methods Arakelyan et al. (2021); Zhu et al. (2022); Bai et al. (2023c); Yin et al. (2023b) incorporate link predictors and search over the symbolic space. In terms of the knowledge domain, existing benchmarks Ren et al. (2020); Ren and Leskovec (2020); Yin et al. (2023a) primarily utilize knowledge graphs containing general factual knowledge, while some studies have focused on commonsense knowledge Fang et al. (2024) and eventuality knowledge Bai et al. (2023a). Research has also shown that such queries can be extended to natural language settings with template or LLM-based approaches Zheng et al. (2024a, b); Zong et al. (2024), enabling broader applications.
@@ 8 Conclusion
In this paper, we studied the first-order logical reasoning capability of transformers with parameterized knowledge under an out-of-distribution generalization setting. We established connections between distribution shifts and logical entailment and comprehensively evaluated transformers on a new knowledge graph query answering benchmark. Detailed results provided empirical evidence of optimal settings throughout the entire pipeline, highlighting the importance of formal language and relative positional encoding. Our results also further motivate the LogiRPE, a logic-aware and more effective self-attention component for first-order reasoning. All our findings give rise to the TEGA, a novel architecture that substantially improves performance and generalizability.
@@
Our study focused on the transformer encoder architecture, which is intuitively more suitable for our task as it generates outputs without the dependency on and interference from the auto-regressive mask found in transformer decoders. Therefore, we did not compare decoder-only and encoder-decoder architectures in our experiments. Additionally, we limited the scope of logical queries to EFO queries with a single free variable, leaving queries with multiple free variables for future exploration.

### v3

@@ Abstract
Transformers, as the fundamental deep learning architecture, have demonstrated great capability in reasoning. This paper studies the generalizable first-order logical reasoning ability of transformers with their parameterized knowledge and how to improve it. Transformers’ capability of first-order reasoning is further captured by whether they can conduct first-order logical entailment, which is quantitatively measured by their performance in answering knowledge graph queries. We establish the connections between (1) two types of distribution shifts studied in out-of-distribution generalization and (2) unseen knowledge and query settings discussed in the task of knowledge graph query answering, which makes it possible to characterize the fine-grained generalizability. Results on our comprehensive dataset showed that transformers outperform previous methods designed particularly for this task and provided detailed empirical evidence about the impact of the input query syntax, token embedding, and transformer architectures on their reasoning capability. Interestingly, our results revealed the mismatch of positional encoding and other design choices of transformer architectures in previous practices. Motivated by this, we propose TEGA, a logic-aware architecture that significantly improves the performance in generalizable first-order logical entailment.
@@
[…]
@@ 1 Introduction
Interestingly, our results demonstrated the superior performance and generalizability of relative positional encoding (RPE) over traditional absolute positional encoding (APE). However, previous studies have only proposed inductive biases within the APE setting, which we prove ineffective in the RPE setting. To fill this gap, we propose TEGA (Transformer Encoder with Guided Attention), a novel modeling methodology that facilitates effective reasoning with logic-aware guidance in self-attention. Our study shows that TEGA substantially improves the performance and generalizability upon transformers under the RPE setting. (§6)
Overall, this work makes three key technical contributions:
  1. 1.
    Benchmark. We conducted comprehensive benchmarking of transformer architectures against baseline models for knowledge graph query answering, using a new dataset that incorporates two distribution shifts.
  2. 2.
    Exploration. We systematically investigated critical design choices throughout the modeling process, including query syntax and permutation, token embedding, and transformer architecture designs.
  3. 3.
    Method. We introduced TEGA, a novel architecture that improves logical entailment performance and generalizability through two inductive biases. We validated its effectiveness through extensive experiments and ablations.
Our code and data can be found at https://github.com/HKUST-KnowComp/TEGA.
@@ 2.2 Knowledge Graph Query Answering
3 Distribution Shifts in KGQA
@@ 4 Evaluation Benchmark
Trans.+Absolute PE Transformer 68.8 18.9 25.8 8.3 43.8 12.7
Trans.+Disentangled PE Transformer 73.6 19.3 29.7 8.7 48.1 13.1
@@
[…]
@@ 4.3 General Benchmarking Results
We evaluate transformer encoders [^2] Decoder-only and Encoder-Decoder yield comparable performance as illustrated in Appendix H. against five existing methods for KG query answering, with details provided in Appendix B. Meanwhile, we compare four positional encoding settings in transformers: Absolute PE is the default sinusoidal positional encoding introduced in the original transformer Vaswani et al. (2023). Relative PE Shaw et al. (2018) applied a learnable matrix based on the relative position of tokens. Disentangled PE He et al. (2021) uses separate vectors to encode the content and position of each token, decoupling the contextual and positional information during self-attention. Rotary PE Su et al. (2023) encodes position information by rotating query/key vectors in the complex space, capturing relative positions through rotation angles.
The experimental results are presented in Table 2. Transformers marginally outperform all baselines in both ID (K) and OOD (K) in seen query types. Both APE and DPE transformers have poor OOD generalizability in the query type dimension. In contrast, RPE and RoPE transformers demonstrates much stronger generalizability, as it can better capture structural information from the distances between tokens and flexibly handle query sequences of unseen lengths. Overall, RPE transformers achieve the best performance and OOD generalizability on first-order logical entailment.
@@ 5 Transformer Architectures and First-order Reasoning Capability
In this section, we further investigate the dependencies between transformers’ performance and multiple design choices throughout the modeling process. Figure 2 depicts the three-stage pipeline for deriving the answer set $\mathcal{A}$ (output $Y$) of query $q$ (input $X$) using transformers. We present the impact of query syntax in Section 5.1, token embeddings in Section 5.2, and transformer architecture in Section 5.3. We summarize the findings in Section 5.4.
@@ 5.1 Study on Query Syntax
Logical queries with the same semantics can be represented in different syntaxes with different formal languages. To investigate the impact of difference in formal languages, we conducted experiments on two formal languages: Lisp-like Syntax (Wang et al., 2021) represents query with fully parenthesized nested formula; EFO Syntax Yin et al. (2023b) represents query with all one-hop atomics connected with conjunctions and disjunctions in parallel. As the Lisp-like syntax has a limited representation scope, we used a subset of our benchmark that includes 13 seen query types and 12 unseen query types that are supported by both languages. The experimental results on the impact of formal language are presented in Table 3. In APE Transformers, there is no difference between the results of the two formal languages. However, with RPE, the EFO syntax achieves much better generalizability in the query type domain. This phenomenon can be explained by the differences in the structural features of the two languages: with a parallel structure, the distances (i.e., relative positions) between tokens with the same logical relationship are more consistent than those in a nested structure. As a result, RPE can better learn the logical relationship between tokens and better generalize to unseen query types.
@@
Furthermore, even for the same query in EFO Syntax, there can be various permutations of one-hop atomics, with the total number of potential permutations growing factorially with the length of the query. Therefore, it is crucial for transformers to robustly handle queries with different permutations. To evaluate this property, we selected five query types (2p, 3p, ip, pi, 2in) and reversed their permutations while maintaining the same semantics in the testing data (see Appendix C). The experimental results on query permutation are presented in Table 4. We observed a significant performance drop after reversing the query in APE transformers. However, for RPE transformers, the performance before and after reversing the permutation remains consistent, demonstrating their robustness in handling queries with modified permutations.
5.2 Study on Token Embeddings
Some methods in KG query answering Arakelyan et al. (2021); Wang et al. (2023b); Xu et al. (2023) utilized pre-trained KG embeddings from link predictors in their query encoding process, while other methods Ren and Leskovec (2020); Zhang et al. (2021); Bai et al. (2023b) randomly initialize the embedding matrix and jointly learn the KG embedding and model parameters from the training process. It’s still unclear that to what extent can pre-trained KG semantics can assist the performance of query answering models. To rigorously assess the impact of pre-trained semantics in token embeddings, we have selected three widely recognized pre-trained KG embeddings: TransE Bordes et al. (2013) models relationships by interpreting them as translations within the embedding space, DistMult Yang et al. (2015) employs a bilinear model wherein entities are represented as vectors and relations as diagonal matrices, ComplEx Trouillon et al. (2016) extends this framework into a complex vector space, allowing for more nuanced inter-entity interactions. In our experiment, all three KG embeddings are frozen during the training stage, in comparison to a baseline with randomly initialized and jointly learned embeddings.
The experimental results are presented in Table 5. Two stronger KGEs, ComplEx and DistMult, can effectively improve the performance of transformer. In contrast, the performance of TransE consistently lags behind the baseline across all three datasets. This discrepancy can be attributed to the baseline model’s embedding matrix learning process, which can be viewed as implicitly training a link predictor using a transformer architecture (e.g., KG-BERT Yao et al. (2019)). This modern approach may naturally outperform earlier models like TransE.
Dataset KGE ID (Q) OOD (Q)
ID (K) OOD (K) ID (K) OOD (K)
FB15k Random 48.1 32.3 35.4 21.5
TransE 44.3 30.3 34.2 20.8
DistMult 49.7 32.6 37.5 21.8
@@ 5.2 Study on Token Embeddings
FB15k-237 Random 54.8 20.0 37.7 15.8
TransE 54.5 20.0 35.0 14.9
@@
NELL995 Random 68.0 18.7 49.8 13.4
TransE 68.5 18.6 48.7 12.5
DistMult 72.2 19.9 51.2 12.9
@@ 5.3 Study on Transformer Architecture
Some existing literature Kotnis et al. (2021); Liu et al. (2022); Xu et al. (2023) has explored the use and design of transformers in KG query answering, in which several inductive biases in transformer architecture were proposed. However, as all these methods focused on different aspects and were evaluated on different benchmarks, a clearer investigation is needed. We implemented two designs from previous works on top of transformers: Adjacency Matrices Masking: an inductive bias proposed in kgTransformer Liu et al. (2022), which aims to enhance reasoning performance by limiting the range of self-attention to one-hop neighboring nodes in the query graph. Directed Distance Encoding: another inductive bias, introduced in Query2Triple Xu et al. (2023), that facilitates information aggregation by injecting a directed distance message into the self-attention mechanism. The experimental results on existing transformer designs are presented in Table 6. Both inductive biases effectively improve generalizability in the query type dimension when using APE. In contrast, under RPE, these improvements are not observed, and the use of adjacency matrices masking even results in a significant decrease in performance. This decrease is attributed to the masking in self-attention, which restricts message passing to only nodes within one hop, thereby limiting the model’s ability to perform complex logical reasoning. Regarding directed distance encoding, the information about directed distances is mostly contained in the relative positions of the linearized sequence, thus the improvements seen with original APE are not sustained.
@@ 5.4 Summary of Empirical Findings
In terms of query syntax, we discovered that employing formal language with parallel structure can significantly enhance OOD generalizability in the query type dimension for RPE transformers. Moreover, we observed that transformers with relative PE can robustly handle queries with different permutations. Regarding transformer architecture, we found that the inductive biases proposed in previous approaches are only effective under APE, but fail to improve upon RPE transformers, which we have shown to be the preferred setting. This mismatch between transformer design and positional encoding settings motivated us to design a methodology that effectively boosts the performance and generalizability of RPE transformers in first-order logical entailment.
@@ 7 Related Work
Transformers have exhibited remarkable performance in various forms of logical reasoning. In natural language inference (NLI) tasks Bowman et al. (2015); Williams et al. (2018), transformers analyze the logical relationship between a premise and a hypothesis, employing deductive reasoning over statements. Furthermore, transformers have been employed for inductive reasoning within rule-based systems Clark et al. (2020). Additionally, transformers have been applied to perform abductive reasoning in both natural language Bhagavatula et al. (2020) and formal language Bai et al. (2024); Gao et al. (2025) settings. Meanwhile, various techniques have been proposed to enhance logical reasoning in language understanding Chen (2023); Pan et al. (2023). Recent studies show that large language models (LLMs) exhibit enhanced logical reasoning abilities Parmar et al. (2024); Li et al. (2025); Fan et al. (2025), even without intermediate reasoning steps Zheng et al. (2025a). Logical entailment also plays a key role in hypothesis verification of LLMs, in areas such as commonsense reasoning Zhao et al. (2024), analogical reasoning Zheng et al. (2025b), and even scientific discovery Takagi et al. (2023); Zheng et al. (2025c).
@@
Prior to this work, extensive research has been conducted on KGQA with parameterized knowledge. Regarding modeling approaches, iterative neural query encoders Ren and Leskovec (2020); Chen et al. (2022); Tsang et al. (2025) design representations for entity sets and execute logical operators iteratively, following the computational graph. In addition, neural-symbolic methods Zhu et al. (2022); Bai et al. (2023c); Yin et al. (2023b) incorporate link predictors and search over the symbolic space. In terms of the knowledge domain, existing benchmarks Ren et al. (2020); Ren and Leskovec (2020); Yin et al. (2023a) primarily utilize knowledge graphs containing general factual knowledge, while some studies have focused on commonsense knowledge Fang et al. (2024) and eventuality knowledge Bai et al. (2023a). Such queries can be extended to natural language settings with template or LLM-based approaches Zheng et al. (2024, 2025d); Zong et al. (2025); Bai et al. (2025a). The setting of KGQA with parameterized knowledge has also been extended to the concept of Neural Graph Databases Ren et al. (2023); Bai et al. (2025b).
@@ 8 Conclusion
This paper investigates the generalizable first-order logical reasoning capabilities of transformers when knowledge is parameterized within model weights. By establishing connections between distribution shifts and logical entailment, we develop a principled framework for evaluating transformer performance on knowledge graph query answering tasks. Our comprehensive benchmark reveals that transformers can effectively perform first-order logical entailment, outperforming methods specifically designed for this task. Through systematic analysis of the entire modeling pipeline, we identify critical design principles that enhance reasoning capabilities. Notably, our findings reveal a mismatch between existing inductive biases and optimal positional encoding strategies, motivating the development of TEGA. This architecture introduces logic-aware self-attention mechanisms that substantially improve both performance and generalizability. Our work demonstrates that careful architectural choices, informed by the logical structure of reasoning tasks, can significantly enhance transformers’ logical entailment ability.
@@
The scope of complex logical queries was limited to EFO queries with a single free variable, with queries containing multiple free variables reserved for future investigation. While LogiRPE incorporates fine-grained logic awareness, it remains dependent on explicit labeling. Future research should explore more generalizable approaches to logic-aware labeling that can accommodate previously unseen logical operators. With respect to knowledge graph scale, the largest knowledge graph employed in our experiments was FB15k, which is considered medium-sized within the knowledge graph literature. Consequently, the scalability of the proposed approaches to large-scale knowledge graphs remains an open question for future investigation.
@@
Acknowledgements
We thank all the anonymous reviewers and meta reviewers for their valuable comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China.